# Robustness is Important: Limitations of LLMs for Data Fitting

### v2

@@ 4 LLMs Have Poor Prediction Robustness with Respect to Task-Irrelevant Variations
Achieving advantageous predictive performance is not the only requirement for data-fitting techniques in practice. We now evaluate LLMs’ prediction sensitivity – whether (and how much) LLMs’ predictions for the same data instances would vary in the presence of task-irrelevant variations to input data, i.e., when we apply changes that do not alter the underlying prediction task in any meaningful way. More specifically, we consider five types of task-irrelevant variations, described as follows:
@@
  • Variable Order: randomly shuffle the order of variables (but maintain the same variable order for each data instance);
  • Number of Digits after decimal point: round each numerical value from 15 digits (the default precision of a double float number in Python) to 10 digits. [^6] While doing so technically reduces the precision of input variables, the applied changes are highly negligible as compared to how we measure and report predictive performance (4 digits after decimal point). We therefore still consider it as a task-irrelevant change.
@@
All task-irrelevant variations lead to nontrivial sensitivity in LLM predictions, which can be as high as 82% under ICL and 24% under SFT. Worse yet, we do not observe any clear patterns across different settings – SFT does not necessarily lead to more robust predictions than ICL; ICL with more in-context examples does not lead to greater robustness; and sensitivity also varies with the distribution of input variables. In other words, LLM-based data fitting and predictions not only suffer from poor robustness with respect to fundamentally task-irrelevant variations, the degree of sensitivity is also hard to anticipate a priori.
@@ 4.1 Prediction Sensitivity Cannot Be Attributed to Randomness
Even though we set LLM generation temperature to 0 in all of the experiments, re-running the exact same prompt may still generate slightly different predictions, due to ties in predicted logit values of candidate tokens. [^7] For instance, if two candidate tokens (e.g., “123” and “124”) share the same predicted logit value during one inference step, some LLMs would randomly choose one of these two candidates as the output token. OpenAI API does not offer precise control over this randomness. We perform statistical tests to ascertain that the prediction sensitivity we observe is not attributed to such randomness in token generation.
@@
If a task-irrelevant change leads to prediction sensitivity that cannot be simply explained away by token generation randomness, we would expect $\Delta_{change}$ to significantly exceed $\Delta_{rand}$ on average. We check this by performing one-sided paired $t$-tests comparing the two quantities across 1,000 testing instances. We find that the differences are statistically significant at 0.001 level for all but two settings (where the prediction sensitivity is small); the only exceptions are SFT with normally distributed variables under variable name change ($p=0.036$, significant at 0.05 level) and under number of digits change ($p=0.064$, significant at 0.1 level).
@@ 4.2 Prediction Sensitivity Cannot Be Avoided by Prompting
We further demonstrate that prediction sensitivity cannot simply be prevented by telling the LLM not to be sensitive with respect to task-irrelevant variations. Let’s consider the Row Order variation as an illustrative example. Under the ICL setting (where training examples are part of the prompt), we insert an extra instruction after the ICL examples that specifies “The presented $K$ examples above are not in any particular order.” This is intended to signal to the LLM that the specific ordering of training examples has no bearing on the data-fitting task.
@@
In general, prompt engineering alone cannot resolve LLMs’ prediction sensitivity with respect to task-irrelevant variations. Even if one can somehow anticipate the variations to guard against ahead of time and insert corresponding prompt instructions, doing so can end up increasing (rather than decreasing) prediction sensitivity.
@@ 5.1 Results
Table 4 summarizes both the predictive performance and the prediction sensitivity of Llama-3-8B-instruct. Overall, it achieves even better predictive performance than GPT-4o-mini under the ICL settings. However, like GPT-4o-mini, its predictions are sensitive with respect to task-irrelevant variations. Note that we set the Llama model to use greedy decoding without random token sampling, so re-running the same setting yields exactly the same set of predictions, thus the observed prediction sensitivity still cannot be attributed to token generation randomness.
@@ 5.2 Attention-Based Mechanism Exploration
We observe similar “U-shaped” attention distributions under the ICL20 setting (with 20 ICL examples that are different from those in ICL10). This further indicates that the “U-shaped” pattern is associated with specific positions in the ICL prompt rather than the training examples we use. See Appendix C for the detailed results.
The uneven attention distributions represent a plausible explanation for LLMs’ prediction sensitivity, because training examples and variables that happen to occupy the beginning / end positions in a prompt end up having larger influence in determining the output tokens. To illustrate this, we repeat the ICL10 experiment but shuffle the variable order. The “U-shaped” attention distributions persist, and variables that show up at beginning / end of each example by random chance now receive greater attentions from the output tokens. We include the results in Appendix C.
@@ 6 Specialized Tabular Foundation Models Are Not Immune to Prediction Sensitivity
Next, we extend our analyses to tabular foundation models specifically designed for data-fitting and prediction tasks. TabPFN Hollmann et al., (2025) represents a state-of-the-art tabular foundation model that has outperformed supervised learning techniques on a wide variety of benchmarking datasets. Importantly, TabPFN recognizes that data fitting should be invariant with respect to variable order and row order, and explicitly tries to facilitate this by implementing a “two-way attention” mechanism where each tabular cell attends to other features in the same row as well as to other values under the same column.
TabPFN can carry out ICL with up to 500 input variables and 10,000 data instances. Therefore, in addition to the ICL10 / ICL20 / ICL500 settings, we also put the entire training dataset (4,000 training instances, denoted as ICL4000) to TabPFN and obtain predictions on the same 1,000 testing instances. Note that TabPFN does not take variable names and requires a standard dataframe format, so we focus on the other three task-irrelevant variations, namely “Variable Order”, “Number of Digits”, and “Row Order”. The computational efficiency of TabPFN allows us to repeat the evaluations 100 times, each with different training / testing datasets simulated from the same data-generation process. We report the mean and standard deviation of performance level and prediction sensitivity in Table 5 for exponentially distributed variables (normally distributed variables yield similar results).
@@
TabPNF achieves impressive predictive performance – with all 4,000 training examples, its average MAE clearly outperforms general-purpose LLMs (GPT-4o-mini and Llama-3-8B-instruct) under both ICL and SFT. However, it is still not immune to task-irrelevant variations, with average prediction sensitivity ranging from 0.20% to 3.04%. Like the Llama model, re-running TabPFN with the same prompt yields exactly the same set of predictions, so the observed prediction sensitivity again cannot be attributed to token generation randomness.
Although the degree of prediction sensitivity may seem small, we stress that task-irrelevant variations should not cause any prediction changes to begin with. As mentioned earlier, this could be viewed as essentially analogous to having a calculator that produces different results when adding up the same set of numbers in a different order. The fact that TabPFN is not immune to task-irrelevant variations points to a deeper issue. Despite TabPFN’s explicit architectural choices to “build in” the invariance with respect to variable order and row order, it cannot fully achieve the said invariance. In contrast, supervised learning techniques are completely invariant to these changes by design. We argue that this highlights a fundamental challenge with using current LLMs for data fitting: although robustness with respect to some task-irrelevant variations could be facilitated to a limited extent with special-purpose tabular foundation models, it generally cannot be guaranteed.
@@ 7 Conclusions
Such prediction sensitivity persists across different learning methods (in-context learning or supervised fine-tuning) as well as different LLMs (general-purpose close- or open-weight LLMs and special-purpose tabular foundation models), and cannot simply be attributed to randomness in token generation. Instead, we provide descriptive evidence that the output tokens are paying more attention to training examples and variable names / values that happen to occupy the beginning or end positions of a prompt, even if those positions are not more important a priori. The non-uniform attention distributions provide a plausible explanation behind spurious sensitivity in LLMs’ predictions.
We believe the absence of even a basic level of prediction robustness brings several challenges. First, it introduces additional complications to the (already difficult) challenge of LLM interpretability (Nanda et al.,, 2023; Conmy et al.,, 2023). Why an LLM produces a certain prediction depends not only on the underlying data but also on specific, task-irrelevant word choices in the prompt. Second, it raises safety, reliability, and trust concerns when LLM-based data fitting and prediction systems are deployed in practice. For instance, if predictions on the same data can be easily altered by changing some variable names, how much would we want to trust those predictions, let alone to make (potentially high-stakes) decisions based on them? Last but not least, people familiar with tabular supervised learning techniques typically take for granted that task-irrelevant variations would not affect their models or predictions. The idiosyncratic choices of LLM users related to data representation can greatly change the results of the same analyses. As one of the consequences, it can further worsen the reproducibility crisis in science (Baker,, 2016; Youyou et al.,, 2023) due to the ever increasing demand for data analysis, data fitting, and predictive modeling in scientific inquiry and the increasing use of LLMs for this purpose.

### v3

@@ 4 LLMs Have Poor Prediction Robustness with Respect to Task-Irrelevant Variations
Achieving advantageous predictive performance is not the only requirement for data-fitting techniques in practice. We now evaluate LLMs’ task-irrelevant prediction sensitivity – whether (and how much) LLMs’ predictions for the same data instances would vary in the presence of task-irrelevant variations to input data, i.e., when we apply changes that do not alter the underlying prediction task in any meaningful way. More specifically, we consider five types of task-irrelevant variations, described as follows:
@@
  • Variable Order: randomly shuffle the order of variables (maintaining the same variable order across data instances);
  • Number of Digits after decimal point: round each numerical value from 15 digits (the default precision of a double float number in Python) to 10 digits. [^6] While doing so technically reduces the precision of input variables, the applied changes are highly negligible as compared to how we measure and report predictive performance (4 digits after decimal point). We therefore still consider it as a task-irrelevant variation.
@@
All task-irrelevant variations lead to nontrivial sensitivity in LLM predictions, which can be as high as 82% under ICL and 24% under SFT. Worse yet, we do not observe any clear patterns across different settings – SFT does not necessarily lead to more robust predictions than ICL; ICL with more in-context examples does not lead to greater robustness; and sensitivity also varies with the statistical distribution of input variables. In other words, LLM-based data fitting and predictions not only suffer from poor robustness with respect to fundamentally task-irrelevant variations, the degree of sensitivity is also hard to anticipate a priori.
@@ 4.1 Prediction Sensitivity Cannot Be Attributed to Randomness
Data fitting and prediction algorithms often involves some randomness. Among tabular supervised learning techniques, random forest is a representative example. A random forest consists of individual decision trees built on randomly selected training instances and feature sets. As a result, applying task-irrelevant variations to the data (e.g., changing the order of variables or rows) may also alter the trained model and its predictions. This gives rise to a natural question: does a random forest exhibit similar degrees of task-irrelevant prediction sensitivity as LLMs? We evaluate this by applying the “Variable Order”, “Number of Digits”, and “Row Order” variations [^7] The “Variable Names” and “Format” variations are not applicable in a tabular data fitting context, and therefore would not cause any prediction sensitivity by definition. and then report the average prediction sensitivity of a random forest model across 100 repetitions (each with different datasets simulated from the same data-generation process). Under exponentially distributed covariates, we find that the prediction sensitivity is 0.01% under “Variable Order” variation, 0.01% under “Number of Digits” variation, and 0.15% under “Row Order” variation, all of which are one to two orders of magnitude smaller than GPT-4o-mini’s corresponding prediction sensitivity. The results under normally distributed covariates are qualitatively consistent. Furthermore, we re-train a random forest model on the exact same data (without applying any task-irrelevant variation) but set a different random seed. We find that average prediction sensitivity caused by varying random seed is 0.14%, which is not significantly different from the sensitivity associated with applying “Row Order” variation ($p=0.66$) and significantly greater than the sensitivity associated with applying “Variable Order” or “Number of Digits” variations ($p<0.001$). In other words, a random forest’s task-irrelevant prediction sensitivity does not exceed what is attributable to randomness in its data-fitting process.
Besides random forest, MLP is another technique that can exhibit task-irrelevant prediction sensitivity. We repeat the above analysis with MLP and find the prediction sensitivity to be 0.01% under “Variable Order” variation, which is not significantly different from sensitivity due to simply varying random seed (0.01%, $p=0.99$). Prediction sensitivity under “Number of Digits” and “Row Order” variations are negligible ($1.38\times 10^{-6}$% and $1.30\times 10^{-10}$% respectively) and significantly lower than the random seed-induced sensitivity ($p<0.001$).
In stark contrast, we find that LLMs’ task-irrelevant prediction sensitivity exceeds what can be explained by randomness. LLM-based data fitting also involves randomness, primarily manifested as randomness in token generation. Moreover, such randomness is not as easy to control as in random forests – even though we set LLM generation temperature to 0 in all our experiments, re-running the exact same prompt may still generate slightly different predictions, due to ties in predicted logit values of candidate tokens and the potential randomness in the LLM’s tie-breaking approach. [^8] For instance, if two candidate tokens (e.g., “123” and “124”) share the same predicted logit value during one inference step, some LLMs would randomly choose one of these two candidates as the output token. OpenAI API does not offer precise control over this randomness. Next, we perform statistical tests to ascertain that the prediction sensitivity we observe cannot be attributed to such randomness in token generation.
@@
If a task-irrelevant change leads to prediction sensitivity that cannot be simply explained away by token generation randomness, we would expect $\Delta_{change}$ to significantly exceed $\Delta_{rand}$ on average. We check this by performing one-sided paired $t$-tests comparing the two quantities across 1,000 testing instances. We find that the differences are statistically significant at 0.001 level for all but two settings (where the prediction sensitivity is small); the only exceptions are SFT with normally distributed variables under variable name change and number of digits change ($p=0.059$ and $p=0.091$, both significant at 0.1 level).
@@ 4.2 Prediction Sensitivity Cannot Be Avoided by Prompting
We further demonstrate that prediction sensitivity cannot be prevented by simply telling the LLM not to be sensitive with respect to task-irrelevant variations. Let’s consider the Row Order variation as an illustrative example. Under the ICL setting (where training examples are part of the prompt), we insert an extra instruction after the ICL examples that states “The $K$ examples presented above are not in any particular order.” This is intended to signal to the LLM that the specific ordering of training examples has no bearing on the data-fitting task.
@@
In general, prompt engineering alone cannot resolve LLMs’ prediction sensitivity with respect to task-irrelevant variations. Even if one can somehow anticipate the variations to guard against ahead of time and insert corresponding prompt instructions, doing so can end up increasing (rather than decreasing) task-irrelevant prediction sensitivity.
@@ 5.1 Results
Table 4 summarizes both the predictive performance and the prediction sensitivity of Llama-3-8B-instruct. Overall, it achieves even better predictive performance than GPT-4o-mini under the ICL settings. However, like GPT-4o-mini, its predictions are sensitive with respect to task-irrelevant variations. Note that we set the Llama model to use greedy decoding without random token sampling, so re-running the same setting yields exactly the same set of predictions; thus, the observed prediction sensitivity still cannot be attributed to token generation randomness. Also, the degrees of prediction sensitivity are much higher than that of random forest or MLP, further indicating that Llama’s task-irrelevant prediction sensitivity is not just driven by randomness.
@@ 5.2 Attention-Based Mechanism Exploration
We observe similar “U-shaped” attention distributions under the ICL20 setting (with 20 ICL examples that are different from those in ICL10). This further indicates that the “U-shaped” pattern is associated with specific positions in the ICL prompt rather than the training examples we use. See SI Appendix, “Additional Analyses of Attention Distributions” Section for the detailed results.
The uneven attention distributions represent a plausible explanation for LLMs’ prediction sensitivity, because training examples and variables that happen to occupy the beginning / end positions in a prompt end up having larger influence in determining the output tokens. To illustrate this, we repeat the ICL10 experiment but shuffle the variable order. The “U-shaped” attention distributions persist, and variables that show up at beginning / end of each example by random chance now receive greater attentions from the output tokens. We include the results in SI Appendix, “Additional Analyses of Attention Distributions” Section.
@@ 6 Specialized Tabular Foundation Models Are Not Immune to Prediction Sensitivity
Next, we extend our analyses to tabular foundation models specifically designed for data-fitting and prediction tasks. TabPFN Hollmann et al., (2025) represents a state-of-the-art tabular foundation model that has outperformed supervised learning techniques on a wide variety of benchmarking datasets. Importantly, the design of TabPFN recognizes that data fitting should be invariant with respect to variable order and row order, and explicitly tries to facilitate this by implementing a “two-way attention” mechanism where each tabular cell attends to other features in the same row as well as to other values under the same column.
TabPFN can carry out ICL with up to 500 input variables and 10,000 data instances. Therefore, in addition to the ICL10 / ICL20 / ICL500 settings, we also put the entire training dataset (4,000 training instances, denoted as ICL4000) to TabPFN and obtain predictions on the same 1,000 testing instances. Note that TabPFN does not take variable names and requires a standard dataframe format, so we focus on the other three task-irrelevant variations, namely “Variable Order”, “Number of Digits”, and “Row Order”. We use the official TabPFNRegressor implementation with default parameters. The computational efficiency of TabPFN allows us to repeat the evaluations 100 times, each with different training / testing datasets simulated from the same data-generation process. We report the mean and standard deviation of performance level and prediction sensitivity in Table 5 for exponentially distributed variables (normally distributed variables yield similar results).
@@
TabPNF achieves impressive predictive performance – with all 4,000 training examples, its average MAE clearly outperforms general-purpose LLMs (GPT-4o-mini and Llama-3-8B-instruct) under both ICL and SFT. Nevertheless, it is not immune to task-irrelevant variations, with average prediction sensitivity ranging from 0.20% to 3.04%, which is still higher than that of canonical tabular machine learning techniques, including random forest or MLP. Moreover, like the Llama model, re-running TabPFN with the same prompt yields exactly the same set of predictions, so the observed prediction sensitivity again cannot be attributed to token generation randomness.
Based on the above results, one might be tempted to conclude that the degree of prediction sensitivity under TabPFN is relatively small, i.e., perhaps not substantial enough to warrant further concern. We argue, however, that this interpretation misses the central issue – that task-irrelevant variations should not cause any prediction changes (beyond what can be explained by randomness in the data-fitting process) for a tool that possesses basic competence in data-fitting tasks. As mentioned earlier, this could be viewed as essentially analogous to having a calculator that produces different results when adding up the same set of numbers in a different order; even if the differences appear “small”, hardly anyone would be comfortable using it as a competent calculation tool.
The fact that TabPFN is not immune to task-irrelevant variations points to a deeper issue. Despite TabPFN’s explicit architectural choices to facilitate the invariance with respect to variable order and row order, it cannot fully achieve the said invariance. In contrast, tabular supervised learning techniques are, by design, either completely invariant to these changes or exhibit minor sensitivity due to randomness (e.g., in random forest or MLP). We argue that this highlights a fundamental challenge with using current LLMs for data fitting: although robustness with respect to some task-irrelevant variations could be facilitated to a limited extent with special-purpose tabular foundation models, it generally cannot be guaranteed.
@@ 7 Conclusions
Such prediction sensitivity persists across different learning methods (in-context learning or supervised fine-tuning) as well as different LLMs (general-purpose close- or open-weight LLMs and special-purpose tabular foundation models), and cannot simply be attributed to randomness in token generation. Instead, we provide descriptive evidence that the output tokens are paying more attention (a) to training examples that happen to occupy the beginning or end positions of a prompt and (b) to variable names / values that happen to occupy the beginning or end positions within each training example, even if those positions are not more important a priori. The non-uniform attention distributions provide a plausible explanation behind spurious sensitivity in LLMs’ predictions.
We believe the absence of even a basic level of prediction robustness brings several challenges. First, it introduces additional complications to the (already difficult) challenge of LLM interpretability (Nanda et al.,, 2023; Conmy et al.,, 2023). Why an LLM produces a certain prediction depends not only on the underlying data but also on specific, task-irrelevant variations of the prompt. Second, it raises safety, reliability, and trust concerns when LLM-based data fitting and prediction systems are deployed in practice. For instance, if predictions on the same data can be easily altered by changing some variable names, how much would we want to trust those predictions, let alone to make (potentially high-stakes) decisions based on them? Last but not least, people familiar with tabular supervised learning techniques typically take for granted that task-irrelevant variations would not affect their models or predictions beyond what can be explained by randomness (e.g., due to having different random seeds); in contrast, the idiosyncratic choices of LLM users related to data representation can greatly change the results of the same analyses. As one of the consequences, it can further worsen the reproducibility crisis in science (Baker,, 2016; Youyou et al.,, 2023) due to the ever increasing demand for data analysis, data fitting, and predictive modeling in scientific inquiry and the increasing use of LLMs for this purpose.
@@ Appendix A Details about Recent Work that Uses LLMs for Data-Fitting
Akhauri et al., (2025) Regressions, specifically for predicting metric outcomes in large complex systems. In-context learning.